) 8. Let C be the triangle which goes from (0,0) to (0,1) to (2,1) and back to (0,0). Let F(x, y) = < 2x”y–sin x,cos y + xy4No need to solve the integral just set up with bounds.
Answer #1

Gwen f(asy) - (2x3y2 - sinx, cosy + 23 y 47 dp = 9|(28 y? - suri x) + (wy +2% y^); }(dal toyi] = $(21° y2 -smix) dx + (cory +in Egnation of line OB (2,1) y - (0, 1) b = EN (0,0) = (3x²y 4 - 4x3y) dady o no This is the required bounds